OpenStudy (anonymous):

Solve using an appropriate method. 5x -9y = 7, 7y -3x = 5

OpenStudy (amorfide):

5x -9y = 7 7y -3x = 5 we can get the same number of x's by multiplying the second equation by 5 and the first equation by 3 this gives 15x-27y=21 -15x+35y=25 now we can add them together to get rid of the x terms this gives 15x+-15x which =0 -27y+35y=8y 21+25=46 our new equation is 8y=46 y=46/8 sub y into one of the original equations to get x

OpenStudy (anonymous):

Thanks! The method appears to be the substitution method. I had to take a break and prepare a snack for the family. Thanks again:)